the sum of 2 numbers is 20. seven times the smaller is 5 more than twice the larger. find each number
Answer by addingup(3677)

x+y = 20 subtract x on both sides: y = 20-x We'll use this value for y next.
7*x = 2y+5 substitute for y:
7*x = 2(20-x)+5
7x = 40-2x+5
9x = 45
x = 5 <-- this is the small number. The big number
5+y = 20
y = 15 this is the big number
:
Happy learning,
John